OpenStudy (camerondoherty):

An equation is shown below: 7x + y = 5 Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

assume an equation in the form: ax + by = c 7x + y = 5 this system can be represented like: \[\left[\begin{matrix}7 & 1 \\ a & b\end{matrix}\right] \left(\begin{matrix}x \\ y\end{matrix}\right) = \left(\begin{matrix}5 \\ c\end{matrix}\right)\]

OpenStudy (anonymous):

find the determinant of first matrix \[\det \left( \left[\begin{matrix}7 & 1 \\ a & b\end{matrix}\right] \right) = 7b - a = 0\]

OpenStudy (camerondoherty):

o.o I havent learned that yet

OpenStudy (anonymous):

It's cramer's rule

OpenStudy (anonymous):

haven't you seen this?

OpenStudy (camerondoherty):

um i dont think so...

OpenStudy (anonymous):

Ok, there's another way to solve it

OpenStudy (anonymous):

we have this equation: 7x + y = 5 y = -7x + 5 so it's slope is -7

OpenStudy (anonymous):

we want to find another equation so that the system is inconsistent

OpenStudy (anonymous):

grafically, this two equations do not touch each other, i.e. they are parallel not coincident

OpenStudy (anonymous):

now that we know that, another equation that we can choose is in this form: y = -7x + k, where k is any number you want, but 5

OpenStudy (anonymous):

let's choose k = 0

OpenStudy (anonymous):

then the system: 7x + y = 5 7x + y = 0 is a inconsistent system

OpenStudy (anonymous):

got it?

OpenStudy (camerondoherty):

thx i was offline sorry

OpenStudy (anonymous):

lmao math -.- throwing cramers rule at a 9th grader

OpenStudy (camerondoherty):

xD

OpenStudy (anonymous):

hahahah